OpenStudy (fibonaccichick666):

In need of some proof assistance. The function \(f(x)=3x^5- 10x^3+15x+1\) is strictly increasing, and therefore, invertible on R. Find \((f^{-1})'(1)\)

OpenStudy (amistre64):

f' = dy/dx (f^-1)' = 1/(dy/dx) = dx/dy

OpenStudy (fibonaccichick666):

I tried to use the inverse function thm, but I ended up with 1/0... By the inverse function thm. since this qualifies due to being strictly monotone and continuous since it is a polynomial, We can use \[(f^{-1})'(y_0)=\frac{1}{f'(x_0)}\]

OpenStudy (fibonaccichick666):

But my issue is that at x=1 f'(x) =0

OpenStudy (amistre64):

f(x) = 3x^5- 10x^3+15x+1 f' = 15x^4 -30x^2 +15 f'(1) = 30-30 = 0 correct

OpenStudy (amistre64):

as such the equation of the inverse of f(x) at x=1 is the line x=1 since we have a vertical slope

OpenStudy (fibonaccichick666):

but I need to find the value of the derivative of the inverse at that point so I wanted to use this example. I'll type it, but I don't know how

OpenStudy (fibonaccichick666):

Let n be a positive interger and let f(x)=x^n for x>0. Then f is strictly increasing on (0,\infty) and its invers is the nth root function g(y)=y^{1/n}.

OpenStudy (fibonaccichick666):

then it just does the computations for g'

OpenStudy (fibonaccichick666):

So, my issue is, how can I apply this thm.?

OpenStudy (fibonaccichick666):

and I need the value of the derivative of the inverse at x=1

OpenStudy (amistre64):

given y= f(x) = x^n : g(y) = g(f(x)) = x if g = y^(1/n)

OpenStudy (fibonaccichick666):

that is the example, but i thought we might be able to use it somehow

OpenStudy (amistre64):

this simply tells us that the inverse of x^n is x^(1/n)

OpenStudy (amistre64):

our f(x) is not strictly increaseing on the interval tho

OpenStudy (fibonaccichick666):

it is

OpenStudy (amistre64):

f'(x) = 0 when x=1 0 isnot increasing

OpenStudy (fibonaccichick666):

it is, it's called strictly increasing ie \(f(x)\le f(y)\le f(z)\)

OpenStudy (fibonaccichick666):

if we observe alphabetical order

OpenStudy (fibonaccichick666):

with a slope of 0, the function remains the same value, but does not decrease

OpenStudy (amistre64):

yeahyeah ... f(x) is increasing

OpenStudy (amistre64):

x^1/3 is an increasing fucntion yes?

OpenStudy (fibonaccichick666):

yep

OpenStudy (amistre64):

what is the slope at x=0?

OpenStudy (fibonaccichick666):

0

OpenStudy (amistre64):

really?

OpenStudy (fibonaccichick666):

oh wait, 1/3 not 3

OpenStudy (fibonaccichick666):

oops, let me look at it

OpenStudy (fibonaccichick666):

1/3 x^{-2/3} at x=0we get a dne, but lim is 0

OpenStudy (amistre64):

there is a vertical slope at x=0 our critical points of interest are when f'=0 or f' is undefinied

OpenStudy (fibonaccichick666):

well, not technically

OpenStudy (fibonaccichick666):

since the function would have to be piecewise, for it to be a vertical slope

OpenStudy (amistre64):

undefined does not mean a tangent line does not exist, since the most standard form of a line is : Ax + By = C

OpenStudy (fibonaccichick666):

otherwise, it could be horizontal

OpenStudy (fibonaccichick666):

also, I think we may be getting off track, as critical points are not allowed to be used here

OpenStudy (amistre64):

the equation of a tangent line that has a vertical slope is simply x = k for some constant k the slope is not undefined ... it is a vertical slope

OpenStudy (fibonaccichick666):

we are not allowed to use that. It has not been proven

OpenStudy (fibonaccichick666):

Literally all I can use is the definition of derivative and the inverse function thm

OpenStudy (fibonaccichick666):

and the definition of continuous if necessary

OpenStudy (amistre64):

a vertical slope is by definition of derivative ... the left and right limit to +inf they have the same limit

OpenStudy (fibonaccichick666):

we don't have left and right limit

OpenStudy (amistre64):

.... youll have to explain to me why we dont

OpenStudy (fibonaccichick666):

because this is an analysis course and if it has not been explicitly proven we are not allowed to use it

OpenStudy (fibonaccichick666):

which is evil, considering I know the answer I just can't arrive there with the tools I am given

OpenStudy (amistre64):

what is your definition of derivaitve ..

OpenStudy (fibonaccichick666):

\[L:=lim_{x \rightarrow c} \frac{f(x)-f(c)}{x-c}\]

OpenStudy (fibonaccichick666):

where L is then =f'(x)

OpenStudy (fibonaccichick666):

sorry f'(c) not x

OpenStudy (amistre64):

and the limit is defined as when the limit from the right and left are the same .... or what is your definition of a limit?

OpenStudy (fibonaccichick666):

it's the epsilon delta definition

OpenStudy (fibonaccichick666):

or the sequential definition, but we are not allowed to use left and right

OpenStudy (fibonaccichick666):

I've been marked wrong for it before

OpenStudy (amistre64):

let c = x+h \[L:=lim_{x \rightarrow (x+h)} \frac{f(x)-f(x+h)}{x-(x+h)}\] \[L:=lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}\]

OpenStudy (fibonaccichick666):

We are not allowed to go there. Unfortunately since that is the def I am more familiar with using

OpenStudy (amistre64):

write up you delta epsilon efinition for me

OpenStudy (fibonaccichick666):

If you don't know it, no offense, but I doubt you can do this proof. That is essential pre-requisite knowledge for it.

OpenStudy (amistre64):

i ahvent taken a real analysis class perse, but i have a read the old textbooks

OpenStudy (fibonaccichick666):

yea, unfortunately, (and believe me I wishhhhhh this was not the case) I don't think you can get this particular proof

OpenStudy (amistre64):

given e > 0, |x-c| > d something like that

OpenStudy (amistre64):

prolly not :)

OpenStudy (fibonaccichick666):

yea, If for every epsilon greater than zero there exists a delta greater than zero such that the absolute value of x-c is less than delta then the absolute value of f(x)-L is less than epsilon

OpenStudy (amistre64):

the absolute value bars in the definition infer a left right limit ... but yeah, without your course material id be at a loss

OpenStudy (fibonaccichick666):

I tried using left and right limits with continuity and I got yelled at :/

OpenStudy (amistre64):

yell back :)

OpenStudy (fibonaccichick666):

hence my lack of wanting to use them again. Haha, well maybe if I was better at writing rigorous proofs I could

OpenStudy (fibonaccichick666):

I ended up just saying that x=1 is implied by g'(y)=1/0

OpenStudy (amistre64):

f(x) is invertible, f'(1) = 0 ... therefore the inverse is 1/0 which is undefiined, or vertical depends on how your course defines 1/0

OpenStudy (fibonaccichick666):

I can't think of how to prove it so hopefully they will just ignore it

OpenStudy (fibonaccichick666):

The issue is that 1/0 was never defined

OpenStudy (fibonaccichick666):

wanna try another?

OpenStudy (amistre64):

the derivative of f(x) is never negative, so f(x) is increasing on any function by the thrm, f(x) is invertible on this interval ... etc

OpenStudy (amistre64):

*f(x) is increasing on any interval

OpenStudy (fibonaccichick666):

Got it

OpenStudy (fibonaccichick666):

here it is essentially: we have (f^{-1})'f(x)<-I equal y!!!=1/f'(x_0) so when does f(x)=1? at x=0 so our x_0=0 and we get 1/15

OpenStudy (fibonaccichick666):

thought you'd like to know